# Multi-Label Bayesian Active Learning with Inter-Label Relationships

### v2

@@ 3 Correlation-aware Multi-label Active Learning
[…]
@@ 3.1 Preliminaries
[…]
@@ 3.2 Correlation Matrix Construction
Positive correlation matrix: The positive correlation matrix $A$ is constructed based on the label-wise dependence. $A$ is a $K\times K$ two dimensional matrix, where each element $A(m,n)$ quantifies the dependency of the presence of label $m$ on label $n$. This dependency is formally computed using Eq. (9). Specifically, $\sum_{i=1}^{L}N(y_{i}^{m}=+1,y_{i}^{n}=+1)$ refers to the count of labeled instances in which both labels $m$ and $n$ appear simultaneously. $P(y^{m}|y^{n})$ gives the likelihood of label $m$ occurring when label $n$ is present. When $m=n$, $A(m,n)$ equals 1. The value of $A(m,n)$ reflects the probability of one label’s existence conditioned on another.
@@
To effectively model these negative correlations, we construct an updated anti-correlation matrix, $NegA$. Maintaining the same format as the positive correlation matrix $A$, $NegA$ is a $K\times K$ two dimensional matrix, where each element $Neg(m,n)$ quantifies the confidence in the absence of label $m$ given the presence of label $n$, as defined in Eq. (10). Specifically, $\sum_{i=1}^{L}N(y_{i}^{m}=-1,y_{i}^{n}=+1)$ represents the count of instances where labels $m$ and $n$ do not co-occur. This asymmetry allows the matrix to capture the nuanced conditional negative relationships, and provide a new perspective towards the label dependencies.
@@ 3.3.3 Hard-to-learn label sampling
Input: Labeled pool: $L$; Unlabeled pool: $U$; Model: $\Theta^{E}={\theta_{1},...,\theta_{E}}$; Query size: $N$; Per-label query size: $N$; Hard to learn query size: $Z$.
@@
4:  Update correlation matrix $A$ and $NegA$ with $L$ via Eq. (9) and Eq. (10)
@@
6:  Select and add $N$ label-wise samples per label to $U^{*}$
7:  Select and add $N$ negative-correlated samples with $NegA$ to $U^{*}$

### v3

@@ 3 Correlation-aware Multi-label Active Learning
Without loss of generality, suppose $L=\{X,Y\}$, $U=\{X\}$ represent the initial collection of training set and unlabeled data samples, where $|U|\gg|L|$; and $y_{i}\in\{-1,+1\}^{k}$ represents the label of the $i_{th}$ example in the $k$ label space. $K$ denotes the total number of labels in the space. Firstly, our model generates the two-dimensional correlation matrix, including both positive correlation and negative correlation, based on the iteratively updated labeled dataset $L$. Then, given a model parameterized by $\theta\in\Theta$, the probability of label $y$ of a data instance $x$ is $P(y|\theta,x)$. We are able to derive the pseudo label as $y^{*}$ based on $\int_{\theta}P(y|\theta,x)P(\theta)d\theta$, where the integration can be approximated by Monte Carlo via ensemble. Considering the model’s learning capability of different categories of data, our model refines the sampling pool into a more preventative subset based on the pseudo labels. And considering the influence of label correlation in quantifying the informativeness of sample, we propose a variation of the beta scoring rule used in Tan et al. [2024]. Its key idea is introduced in Section 3.4. Finally, the clustering approach assures the diversity in sampling. Fig. 1 illustrates the overall flowchart of our proposed framework.
@@ 3.1 Preliminaries
$\mathcal{L}(1|p)$ $=\mathcal{L}_{1}(1-p)=\textstyle\int^{1}_{p}p^{\alpha-1}(1-p)^{\beta}dp$ (7)
$\mathcal{L}(0|p)$ $=\mathcal{L}_{0}(p)=\textstyle\int^{p}_{0}p^{\alpha}(1-p)^{\beta-1}dp$ (8)
@@ 3.2 Correlation Matrix Construction
Positive correlation matrix: The positive correlation matrix AA is constructed based on the label-wise dependence. AA is a $K\times K$ two dimensional matrix, where each element $A(m,n)$ quantifies the dependency of the presence of label $m$ on label $n$. This dependency is formally computed using Eq. (9). Specifically, $\sum_{i=1}^{L}N(y_{i}^{m}=+1,y_{i}^{n}=+1)$ refers to the count of labeled instances in which both labels $m$ and $n$ appear simultaneously. $P(y^{m}|y^{n})$ gives the likelihood of label $m$ occurring when label $n$ is present. When $m=n$, $A(m,n)$ equals 1. The value of $A(m,n)$ reflects the probability of one label’s existence conditioned on another.
@@
To effectively model these negative correlations, we construct an updated anti-correlation matrix, NegANegA. Maintaining the same format as the positive correlation matrix AA, NegANegA is a $K\times K$ two dimensional matrix, where each element $Neg(m,n)$ quantifies the confidence in the absence of label $m$ given the presence of label $n$, as defined in Eq. (10). Specifically, $\sum_{i=1}^{L}N(y_{i}^{m}=-1,y_{i}^{n}=+1)$ represents the count of instances where labels $m$ and $n$ do not co-occur. This asymmetry allows the matrix to capture the nuanced conditional negative relationships, and provide a new perspective towards the label dependencies.
@@ 3.3.3 Hard-to-learn label sampling
Input: Labeled pool: $L$; Unlabeled pool: $U$; Model: $\Theta^{E}={\theta_{1},...,\theta_{E}}$; Query size: NN; Per-label query size: NN; Hard to learn query size: $Z$.
@@
4: Update correlation matrix AA and NegANegA with $L$ via Eq. (9) and Eq. (10)
@@
6: Select and add NN label-wise samples per label to $U^{*}$
7: Select and add NN negative-correlated samples with NegANegA to $U^{*}$